OpenStudy (anonymous):

How can you represent the terms of a sequence explicitly? How can you represent them recursively? Provide mathematical examples to support your opinions.

OpenStudy (anonymous):

If a sequence has the numbers 1, 5, 9, 13, 17. Explicitly a^n= 4n+5 Recursively a^n = a^(n-1)+4

OpenStudy (anonymous):

Is my answer correct?

OpenStudy (turingtest):

not quite, do you propose to start with n=1 or n=0 ?

OpenStudy (anonymous):

Shouldn't it start with n=1? (I just started studying this)

OpenStudy (turingtest):

you can do it either way but look what happens in your explicit formula when n=1. what is the value?

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

That wouldn't work out if so..

OpenStudy (turingtest):

right, so how can we get 5? we would have to start with n=0 like I say, that's ok, but then how can we get the 1?

OpenStudy (anonymous):

I'm not sure how I would go about that, I would have to write a new explicit formula right?

OpenStudy (turingtest):

yeah, but it won't be too different consider the first terms written differently (let's start from n=0)\[a_0=1\\a_1=1+4\\a_2=1+4+4\\~~~~~\vdots\]any ideas yet?

OpenStudy (anonymous):

a^n=1+4^n

OpenStudy (turingtest):

well ^ usually means exponent...

OpenStudy (turingtest):

do you mean\[a_n=1+4n\]

OpenStudy (anonymous):

Yes sorry. That should've been written like a_n=1+4^n right?

OpenStudy (turingtest):

again, 4^n is exponent do you mean 4 to the power of n?

OpenStudy (anonymous):

Yes, I thought that was how you typed it?

OpenStudy (turingtest):

it is, but the answer should not be an exponent we are adding 4 every time. exponent would be multiplying by 4 every time

OpenStudy (anonymous):

Ohhhhh! I just realized what I was doing. \[a _{n}=1+4n\]

OpenStudy (turingtest):

yes, much better :)

OpenStudy (anonymous):

Okay, so how would we figure out the recursive equation?

OpenStudy (turingtest):

ok so your recursive is almost right, except again you have no way of getting to the first term. however to do this for recursive functions, the method is to simply define the first term, which if you think about it is necessary. so what is the value of \(a_0\) ?

OpenStudy (anonymous):

In my equation the value of my first term would be 3.... But I need to find a way to write it so it is 1.

OpenStudy (turingtest):

oh, ok look you are writing exponents all over the place. there should be no exponents. what you want instead of exponents on the left are subscripts, which you can write with your keyboard as a_n=a_(n-1)+4 (I know it looks ugly, you can use the equation editor or LaTeX if you want to make it like mine)

OpenStudy (turingtest):

you are writing\[a^n=a^{n-1}+4\]it should be\[a_n=a_{n-1}+4\]

OpenStudy (anonymous):

Okay.. I wasn't sure how I was supposed to write subscript.

OpenStudy (turingtest):

well how did you get your first value as 3 ?

OpenStudy (anonymous):

by plugging it into my old recursive equation.

OpenStudy (anonymous):

Not 3... sorry that should be 4.

OpenStudy (turingtest):

what did you plug in? how did you get the value of \(a_{n-1}\)?

OpenStudy (anonymous):

Shouldn't I plug in 1?

OpenStudy (turingtest):

for n? then you get\[a_1=a_0+4\]so you still have the problem of how to get \(a_0\)

OpenStudy (turingtest):

you just plugged in \(0\) for \(a_0\), which is unjustified

OpenStudy (anonymous):

Okay so I'm starting with my first term (a_0) and I'm plugging in 1 for that right?

OpenStudy (turingtest):

yes, so we need to state that in our formula recursive formulas always need a base case like this, otherwise they can never get started

OpenStudy (turingtest):

so to make our recursive formula complete we need to actually say\[a_0=1\]in addition to what you wrote

OpenStudy (turingtest):

does that make any kind of sense? that without a base case where we define one term to be a particular value, our recursive formula won't work?

OpenStudy (anonymous):

Yeah that makes sense.

OpenStudy (turingtest):

Well I hope that helped See you around!

OpenStudy (anonymous):

Wait what was my recursive formula going to be?

OpenStudy (turingtest):

\[a_0=1\\a_n=a_{n-1}+4\]